solve x^2du/dx+y^2du/dy=0


1
Expert's answer
2020-10-12T18:59:42-0400

The subsidiary equations are given by


"\dfrac{dx}{x^2}=\dfrac{dy}{y^2}=\dfrac{du}{0}"

"\dfrac{dx}{x^2}=\dfrac{dy}{y^2}"

"\int\dfrac{dx}{x^2}=\int\dfrac{dy}{y^2}"

"-\dfrac{1}{x}=-\dfrac{1}{y}+C_1"

"\dfrac{1}{y}-\dfrac{1}{x}=C_1"

"\dfrac{dx}{x^2}=\dfrac{du}{0}"

"du=0=>u=C_2"

The desired solution is "f(C_1, C_2)=0" or


"f(\dfrac{1}{y}-\dfrac{1}{x}, u)=0"